The expression for the number of molecules in a Maxwellian gas having speeds in

the range v to v + dv is


dNV=4(3.14)N((M/(2(3.14)kBT))3/2 V2 exp[-(mv2/2kBT)]dV)


Using this relation, obtain an expression for average speed. Also, plot Maxwellian

distribution function versus speed at three different temperatures.


1
Expert's answer
2021-03-08T08:18:55-0500

"dN_v=4\pi N(\frac{m}{2\pi k_bT})^{\frac 32} v^2 e^{-\frac{mv^2}{2k_b T}}dv,"

"F(v)=4\pi (\frac{m}{2\pi k_bT})^{\frac 32} v^2 e^{-\frac{mv^2}{2k_b T}}dv,"


"<v>=\int_0^{\infin}vF(v)dv=4\pi (\frac{m}{2\pi k_bT})^{\frac 32} \int_0^{\infin}v^3 e^{-\frac{mv^2}{2k_b T}}dv=\sqrt{\frac{8k_b T}{\pi m}}"